Suppose that stock of ACME Incorporated has a required return of 10% an initial dividend of...

  1. Suppose that stock of ACME Incorporated has a required return of 10% an initial dividend of $1 and 3% expected growth of its dividend. Find its price using the Gordon Growth Model.
Answer #1

ANSWER:

Price = dividend / (return - growth) = 1 / (10% - 3%) = 1 / 7% = 14.28

so the price is $14.28